OpenStudy (anonymous):

the area of a typical room in this hotel is 325ft . What is the perimeter of the room if this is a typical room? show your work

jhonyy9 (jhonyy9):

what is your opinion how you start it ? what form has a room ?

OpenStudy (anonymous):

what do u mean?

jhonyy9 (jhonyy9):

triangle ?

jhonyy9 (jhonyy9):

how you get a perimeter of a room ?

OpenStudy (anonymous):

it doesnt say

OpenStudy (anonymous):

i think u add?

jhonyy9 (jhonyy9):

so you know it too sure that a typical room has 4 sides , general has a room a rectangle or square form but typical room has rectangle form sure - yes ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so do i add or what do i do

jhonyy9 (jhonyy9):

add what ?

OpenStudy (anonymous):

the perimeter or no

jhonyy9 (jhonyy9):

for perimeter you need adding the length of these 4 sides but you know it sure that from these 4 sides 2 by 2 has equale length - yes ?

jhonyy9 (jhonyy9):

suppose this is a rectangle form room - ok. ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

please collaborate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im really confused

jhonyy9 (jhonyy9):

so for example just now use these letters for perimeter - how you get it ?

jhonyy9 (jhonyy9):

we have clarified above that for getting the perimeter we need adding the sides lebgth - yes ?

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

so than in this case how many will be this perimeter using to calcule it just these letters ?

jhonyy9 (jhonyy9):

P=?

OpenStudy (anonymous):

im thinking so hard right now

jhonyy9 (jhonyy9):

adding 2 long side + 2 width side

OpenStudy (anonymous):

oh ok

jhonyy9 (jhonyy9):

P = ?

OpenStudy (anonymous):

on my paper it says the sides are 5in. and 1 inch equals 2.5 feet

jhonyy9 (jhonyy9):

one side has 2,5 feet length - so with this 2,5 ft multiplied by 4 you not get 325 ft perimeter - yes ?

OpenStudy (anonymous):

yeah i guess

jhonyy9 (jhonyy9):

realy ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

when i multiply i get 10

jhonyy9 (jhonyy9):

what multiplie you to get 10 ?

jhonyy9 (jhonyy9):

so you know that P = 2a+2b

OpenStudy (anonymous):

four is what i multiplied by 2.5 to get 10

jhonyy9 (jhonyy9):

so you know that perimeter of a room equal 2a+2b where a and b are sides length so than you can writing that 325 =2a+2b - yes ?

jhonyy9 (jhonyy9):

you said above that on your paper said that one side has 5 in length and 1 inch = 2,5 feet yes ?

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

than 1 inch = 2,5 feet 5 in = x feet --------------------- x = 5*2,5 = ?

jhonyy9 (jhonyy9):

x=12,5 feet but indifferent because 4*12,5 for perimeter not will give you result 325 ft never - so than what is there on your paper ?

jhonyy9 (jhonyy9):

so you know that P=325 ft and P=2a+2b =2(a+b) 325 = 2(a+b) divide both sides by 2 162,5 =a+b

OpenStudy (anonymous):

it says the area of the typical room is 325ft what is the perimeter of the room

OpenStudy (anonymous):

oh yeah

jhonyy9 (jhonyy9):

ohhhh sorry

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

np

jhonyy9 (jhonyy9):

so than you need getting length time width =325 - yes ?

OpenStudy (anonymous):

yas

jhonyy9 (jhonyy9):

because area of a rectangle is equal length time width

OpenStudy (anonymous):

yup

jhonyy9 (jhonyy9):

so do you have any opinion about these numbers what multiplied result to you 325 ?

jhonyy9 (jhonyy9):

so you know that 20*15=300

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

can you get them ?

jhonyy9 (jhonyy9):

so you need getting 2 numbers what multiplied will result 325

OpenStudy (anonymous):

oh i get it now

jhonyy9 (jhonyy9):

can you writing me here please ?

OpenStudy (anonymous):

20*16 is close it =320

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

thx

jhonyy9 (jhonyy9):

so than you have got them you can calcule easy the perimeter of a room

OpenStudy (anonymous):

how do i give medal

jhonyy9 (jhonyy9):

first we need endong it with perimeter - ok. ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

please write here these sides of the room and calcule the perimeter - how many will be ?

OpenStudy (anonymous):

i have to go

OpenStudy (anonymous):

thx tho

jhonyy9 (jhonyy9):

you have got them - so to give me medal just push on best response in line with my name